# New inequalities for the extended Euler-Poincaré theorem

### v1

@@ Abstract
In [Acta Math. 161 (3–4) (1988) 279–303], Björner and Kalai extended the classic Euler-Poincaré theorem by introducing certain nonlinear relations between the $f$-vector and the Betti sequence of simplicial complexes. In this paper, we present an equivalent characterization of Björner and Kalai’s nonlinear relations using a number-theoretic approach. Moreover, we strengthen a result of Björner and Kalai concerning the maximal element of Betti sequences with respect to a fixed $f$-vector and the minimal element of $f$-vectors with respect to a fixed Betti sequence.
Keywords: Euler-Poincaré theorem, simplicial complex, $f$-vector, Betti number
@@ 1 Introduction
This unique expansion allows us to define the following operators:
@@
$\partial^{k-1}(n)$ $=\binom{a_{k}-1}{k}+\binom{a_{k-1}-1}{k-1}+\cdots+\binom{a_{i}-1}{i},$
$d^{k-1}(n)$ $=\binom{a_{k}+1}{k}+\binom{a_{k-1}+1}{k-1}+\cdots+\binom{a_{i}+1}{i}.$
Furthermore, let $\delta_{k-1}(n)$ denote the number of indices $t$ in the unique expansion (1) for which $a_{t}=t$. For completeness, we define
$\partial_{k-1}(0)=\partial^{k-1}(0)=d^{k-1}(0)=\delta_{k-1}(0)=0.$
Let $\mathbb{N}_{0}^{(\infty)}$ denote the set of ultimately vanishing sequences of nonnegative integers. For a sequence $f=(f_{0},f_{1},\ldots)\in\mathbb{N}_{0}^{(\infty)}$, we define
@@
Let $\Delta$ be a finite simplicial complex. The $f$-vector of $\Delta$ is the sequence $f=(f_{0},f_{1},\ldots)$, where $f_{i}$ counts the number of faces of dimension $i$ in $\Delta$, that is,
$f_{i}=\#\{F\in\Delta\mid\dim F=i\}\quad\text{for }i\geq 0.$
@@
$\beta_{i}=\dim_{\boldsymbol{k}}\widetilde{H}_{i}(\Delta;\boldsymbol{k})\quad\text{for }i\geq 0.$
The well-known Euler–Poincaré theorem (see [6, 7]) states that
$\chi_{-1}=\chi_{-1}(f,\beta)=\sum_{j\geq 0}(-1)^{j}(f_{j}-\beta_{j})=1.$ (2)
The Euler–Poincaré formula (2) establishes a linear relation between the $f$-vector $f$ and the reduced Betti sequence $\beta$ of $\Delta$. A natural question is whether there exist other linear relations between $f$ and $\beta$ beyond (2). However, a result of Mayer [5] shows that no such relations exist.
The Euler–Poincaré theorem was significantly extended by Björner and Kalai [1] in 1988, who discovered $d$ additional nonlinear relations (where $d=\dim\Delta$) connecting $f$ and $\beta$. Together with (2), these relations completely characterize all possible pairs $(f,\beta)$ that can occur as the $f$-vector and Betti sequence of a finite simplicial complex.
Theorem 1.1 (Björner and Kalai, [1, Theorem 1.1]).
@@ 2. $(b)$
The set $\mathbb{N}_{0}^{(\infty)}$ and all its subsets are naturally equipped with the componentwise partial order, defined by
An ordered pair $(f,\beta)$ of sequences in $\mathbb{N}_{0}^{(\infty)}$ is called compatible, denoted by $f\sim\beta$, if there exists a simplicial complex with $f$-vector $f$ and Betti sequence $\beta$. By Theorem 1.1, this relation is purely combinatorial and independent of the field characteristic.
In this paper, we employ a number-theoretic approach to present an equivalent reformulation of Theorem 1.1.
Theorem 1.2.
  1. $(a)$
  2. $(b^{*})$
    $\chi_{-1}=1$, and $\partial_{k}(\chi_{k}+\beta_{k})\leq\chi_{k-1}-\epsilon_{k}$, for all $k\geq 1$,
  3. $(b^{+})$
    $\chi_{-1}=1$, and $0\leq f_{k}-\chi_{k-1}\leq\partial^{k}(f_{k})$, for all $k\geq 1$.
Remark 1.1.
It is worth mentioning that condition $(b^{*})$ in Theorem 1.2 refines condition $(b)$ from Theorem 1.1, while condition $(b^{+})$ establishes a new characterization of compatible pairs $(f,\beta)$. Moreover, for any simplicial complex, we have $f_{k}-\chi_{k-1}=\dim Z_{k}$, where $Z_{k}$ denotes the space of $k$-cycles (see [1, Remark 4.4]). Therefore, the homological interpretation of the inequality $f_{k}-\chi_{k-1}\leq\partial^{k}(f_{k})$ in Theorem 1.2 $(b^{+})$ yields
$\dim Z_{k}\leq\partial^{k}(f_{k}).$
In their seminal work [1], Björner and Kalai further established the following result about compatible pairs $(f,\beta)$.
Theorem 1.3 (Björner and Kalai, [1, Theorem 1.2, Theorem 5.3]).
@@ 6. $(f)$
We now strengthen partial results from Theorem 1.3 by sharpening the inequalities in parts $(c)$ and $(f)$, and providing an equivalent reformation of part $(d)$.
@@ 6. $(f^{*})$
    $\phi(\psi(f)+\delta_{+}(f))=f-\delta(f)$ for all $f$-vectors $f$. Consequently, $\phi(\psi(f))\leq f-\delta(f)-\delta_{+}(f)$, with equality holding if and only if $\delta(f)=0$.
2 Proofs of main results
In this section, we will prove Theorems 1.1 and 1.3. We begin by introducing a result of purely number-theoretic nature.
Theorem 2.1.
Let $f_{k}$, $\chi_{k-1}$, $\chi_{k}$ and $\beta_{k}$ be nonnegative integers satisfying $f_{k}=\chi_{k-1}+\chi_{k}+\beta_{k}$, and let $\epsilon_{k}$ be an integer with $0\leq\epsilon_{k}\leq\delta_{k}(f_{k})$. Then the following statements hold:
@@ 1. $(a)$
    $\partial_{k}(\chi_{k}+\beta_{k})\leq\chi_{k-1}-\epsilon_{k}$ if and only if $0\leq f_{k}-\chi_{k-1}\leq\partial^{k}(f_{k})$;
  2. $(b)$
    $f_{k}-\chi_{k-1}=\partial^{k}(f_{k})$ if and only if $\partial_{k}(\chi_{k}+\beta_{k})+\delta_{k}(f_{k})=\chi_{k-1}$;
  3. $(c)$
    $\partial_{k}(\chi_{k}+\beta_{k})=\chi_{k-1}$ if and only if $f_{k}-\chi_{k-1}=\partial^{k}(f_{k})$ and $\delta_{k}(f_{k})=0$.
@@ 3. $(c)$
such that $a_{k}>a_{k-1}>\cdots>a_{i}\geq i\geq 1$. This expansion is called the $(k-1)$-dimensional representation of $n$.
@@
are the $k$-dimensional representations of $n$ and $m$, respectively. Then $n<m$ if and only if either $i>j$ and $a_{t}=b_{t}$ for all $t\in\{i,\ldots,k+1\}$, or the maximal index $t\in\{\max\{i,j\},\ldots,k+1\}$ with $a_{t}\neq b_{t}$ satisfies $a_{t}<b_{t}$. Moreover, $n=m$ if and only if $i=j$ and $a_{t}=b_{t}$ for all $t\in\{i,\ldots,k+1\}$.
@@
Using Pascal’s identity $\binom{p}{q}=\binom{p-1}{q}+\binom{p-1}{q-1}$ recursively, we obtain
@@
for any integers $10\leq d\leq c+$ 1. This implies that
@@
whenever $c_{p}>d_{p}$ and $d_{p}>d_{p-1}>\cdots>d_{l}\geq l\geq 1$. Let $s=\max\{i,j\}$. If $a_{t}=b_{t}$ for all $t\in\{s,\ldots,k+1\}$, then $n<m$ when $i>j$, $n>m$ when $i<j$, and $n=m$ when $i=j$. If there exists some index $t\in\{s,\ldots,k+1\}$ such that $a_{t}\neq b_{t}$, let $t_{0}=\max\{t\in\{s,\ldots,k+1\}:a_{t}\neq b_{t}\}$. Then it follows from (3) that $n<m$ when $a_{t_{0}}<b_{t_{0}}$, and $n>m$ when $a_{t_{0}}>b_{t_{0}}$. ∎
@@
The case $n=0$ is trivial, so we may assume $n\geq 1$. By Lemma 2.1, the $k$-dimensional representation of $n$ is given by
@@
where $a_{k+1}>a_{k}>\cdots>a_{i}\geq i\geq 1$. Then
@@
The case $n=0$ is trivial, so we may assume $n\geq 1$. By Lemma 2.1, the $k$-dimensional representation of $n$ is given by
@@
where $a_{k+1}>a_{k}>\cdots>a_{i}\geq i\geq 1$. If $a_{t}=t$ for all $t\in\{i,\ldots,k+1\}$, then $\delta_{k}(n)=n=k+2-i$, $\partial^{k}(n)=\partial_{k}(\partial^{k}(n))=0$, and we are done. Now suppose there exists some $t\in\{i,\ldots,k+1\}$ such that $a_{t}\neq t$. Let $t_{0}=\min\{t\in\{i,\ldots,k+1\}:a_{t}\neq t\}$. Then $a_{t}=t$ whenever $t<t_{0}$, and $a_{t}>t$ whenever $t\geq t_{0}$. According to the definition of the function $\delta_{k}$, we obtain
@@
Thus it follows from Lemma 2.3 that
@@
Let $n$ and $m$ be nonnegative integers. Then the following statements hold.
@@
First we consider $(a)$. If $d^{k}(n)=m$, by Lemma 2.2, we obtain $i=j$ and $a_{t}+1=b_{t}$ for all $t\in\{i,\ldots,k+1\}$. Therefore, $n=\partial^{k}(m)$. If $d^{k}(n)<m$, again by Lemma 2.2, either $i>j$ and $a_{t}+1=b_{t}$ for all $t\in\{i,\ldots,k+1\}$, or the maximal index $t\in\{\max\{i,j\},\ldots,k+1\}$ with $a_{t}+1\neq b_{t}$ satisfies $a_{t}+1<b_{t}$. For the former case, we have
@@
$n$ $=\binom{a_{k+1}}{k+1}+\binom{a_{k}}{k}+\cdots+\binom{a_{i}}{i}$ (4)
@@
Now consider $(b)$. If $n=m$, then $\partial_{k}(n)=\partial_{k}(m)$ by definition. Thus we can assume that $n<m$. Again by Lemma 2.2, we have that either $i>j$ and $a_{t}=b_{t}$ for all $i\leq t\leq k+1$, or the maximal index $t\in\{\max\{i,j\},\ldots,k+1\}$ with $a_{t}\neq b_{t}$ satisfies $a_{t}<b_{t}$. For the former case, we have
@@
In the latter case, let $t_{0}=\max\{t\in\{\max\{i,j\},\ldots,k+1\}:a_{t}\neq b_{t}\}$. Then $a_{t_{0}}<b_{t_{0}}$, and we get
@@
If $\partial_{k}(\chi_{k}+\beta_{k})\leq\chi_{k-1}-\epsilon_{k}$, by Lemma 2.3,
$d^{k}(f_{k}-\chi_{k-1})=d^{k}(\chi_{k}+\beta_{k})=\partial_{k}(\chi_{k}+\beta_{k})+\chi_{k}+\beta_{k}\leq\chi_{k-1}-\epsilon_{k}+\chi_{k}+\beta_{k}=f_{k}-\epsilon_{k}.$
@@
$f_{k}-\chi_{k-1}\leq\partial^{k}(f_{k}-\epsilon_{k})=\partial^{k}(f_{k}).$
Conversely, if $f_{k}-\chi_{k-1}\leq\partial^{k}(f_{k})$, by Lemma 2.5 $(b)$, we have
$\partial_{k}(f_{k}-\chi_{k-1})\leq\partial_{k}(\partial^{k}(f_{k})),$
@@
$f_{k}-\chi_{k-1}+\partial_{k}(f_{k}-\chi_{k-1})\leq\partial^{k}(f_{k})+\partial_{k}(\partial^{k}(f_{k})).$
@@
$\partial_{k}(f_{k}-\chi_{k-1})$ $\leq\chi_{k-1}-(f_{k}-\partial^{k}(f_{k})-\partial_{k}(\partial^{k}(f_{k})))=\chi_{k-1}-\delta_{k}(f_{k})\leq\chi_{k-1}-\epsilon_{k}.$
Therefore, we conclude that $\partial_{k}(\chi_{k}+\beta_{k})=\partial_{k}(f_{k}-\chi_{k-1})\leq\chi_{k-1}-\epsilon_{k}$. This proves $(a)$.
If $f_{k}-\chi_{k-1}=\partial^{k}(f_{k})$, then $\partial_{k}(f_{k}-\chi_{k-1})=\partial_{k}(\partial^{k}(f_{k}))$. This implies
$f_{k}-\chi_{k-1}+\partial_{k}(f_{k}-\chi_{k-1})=\partial^{k}(f_{k})+\partial_{k}(\partial^{k}(f_{k})),$ (5)
@@
$\partial_{k}(\chi_{k}+\beta_{k})+\delta_{k}(f_{k})=\chi_{k-1}$
by Lemma 2.4. Conversely, if $\partial_{k}(\chi_{k}+\beta_{k})+\delta_{k}(f_{k})=\chi_{k-1}$, then we recover (5). By Lemma 2.3, it follows that
$d^{k}(f_{k}-\chi_{k-1})=d^{k}(\partial^{k}(f_{k})).$ (6)
@@
Suppose that $f_{k}-\chi_{k-1}=\partial^{k}(f_{k})$. According to $(b)$, we have $\partial_{k}(\chi_{k}+\beta_{k})+\delta_{k}(f_{k})=\chi_{k-1}$. If $\delta_{k}(f_{k})=0$ then $\partial_{k}(\chi_{k}+\beta_{k})=\chi_{k-1}$, and if $\delta_{k}(f_{k})\neq 0$, then $\partial_{k}(\chi_{k}+\beta_{k})\neq\chi_{k-1}$. On the other hand, if $\partial_{k}(\chi_{k}+\beta_{k})=\chi_{k-1}$, by Lemma 2.3,
$d^{k}(f_{k}-\chi_{k-1})=d^{k}(\chi_{k}+\beta_{k})=\partial_{k}(\chi_{k}+\beta_{k})+\chi_{k}+\beta_{k}=\chi_{k-1}+\chi_{k}+\beta_{k}=f_{k}.$
Combining this with Lemma 2.5 $(a)$, we obtain $f_{k}-\chi_{k-1}=\partial^{k}(f_{k})$, and so $(c)$ follows.
@@
By Theorem 1.1 and Theorem 2.1, we obtain the proof of Theorem 1.2 immediately. In order to prove Theorem 1.4, we also need the Kruskal-Katona theorem [3, 4], which gives a complete characterization of $f$-vectors of simplicial complexes.
@@
First observe that $(d^{*})$ follows from Theorem 1.3 and Theorem 2.1 immediately. Now suppose that $f$ is the $f$-vector of some simplicial complex. Let $\beta\in B_{f}$. Since $f\sim\beta$, by Theorem 1.2, we have $\chi_{-1}=1$, and $f_{k}-\chi_{k-1}\leq\partial^{k}(f_{k})$ for all $k\geq 1$. Also note that $f_{0}-\chi_{-1}=f_{0}-1=\partial^{0}(f_{0})$. Hence, for each $k\geq 0$,
@@
Furthermore, according to [1, Theorem 5.3], the upper bound in (7) can be achieved uniformly, that is,
@@
Next we consider $(f^{*})$. According to $(c^{*})$ and $(d^{*})$, if $f$ is an $f$-vector with $\delta(f)=0$, then $\phi(\psi(f))=f$. Now suppose that $f$ is an arbitrary $f$-vector. By Lemma 2.6, it is easy to see that $f-\delta(f)$ is also an $f$-vector. Moreover, by definition, we have $\delta(f-\delta(f))=0$. Thus $\phi(\psi(f-\delta(f)))=f-\delta(f)$ by the above argument. Note that, for any $k\geq 0$,
@@
By [1, Theorem 5.2], the mapping $\phi$ is injective and order-preserving. Therefore, we conclude that
@@
where the equalities hold everywhere if and only if $\delta(f)=0$.

### v3

# A note on the Björner–Kalai theorem

@@ Abstract
In 1988, Björner and Kalai used combinatorial shadow functions to characterize the maximal Betti sequence for a given $f$-vector and the minimal $f$-vector for a given Betti sequence. Their description of the maximal Betti sequence was expressed through a set of inequalities. In this paper, we introduce an error function $\delta_{k}$ associated with the combinatorial shadow functions and use it to sharpen these inequalities into exact equalities. As a corollary, we obtain an equivalent form of Björner and Kalai’s characterization of all possible pairs $(f,\beta)$ that can occur as the $f$-vector and Betti sequence of a simplicial complex. Moreover, combining our results with a previous result of Björner in 2011, we derive a new number-theoretic inequality concerning the count of odd square-free integers with a specified number of prime factors.
Keywords: combinatorial shadow function, simplicial complex, $f$-vector, Betti number
@@ 1 Introduction
This unique expansion allows one to define the following combinatorial shadow functions:
@@
$\partial^{k-1}(n)$ $=\binom{a_{k}-1}{k}+\binom{a_{k-1}-1}{k-1}+\cdots+\binom{a_{i}-1}{i}$.
For completeness, let $\partial_{k-1}(0)=\partial^{k-1}(0)=0$.
Let $\mathbb{N}_{0}^{(\infty)}$ denote the set of ultimately vanishing sequences of nonnegative integers. The set $\mathbb{N}_{0}^{(\infty)}$ and all its subsets are naturally equipped with the componentwise partial order, defined by
@@
Let $\Delta$ be a finite simplicial complex. The $f$-vector of $\Delta$ is the sequence $f=(f_{0},f_{1},\ldots)$, where
$f_{i}=\#\{F\in\Delta\mid\dim F=i\},\quad\text{for each }i\geq 0.$
@@
$\beta_{i}=\dim_{\boldsymbol{k}}\widetilde{H}_{i}(\Delta;\boldsymbol{k}),\quad\text{for each }i\geq 0.$
Using the combinatorial shadow function $\partial_{k}$, Björner and Kalai [2] provided a complete characterization of all possible pairs $(f,\beta)$ that may occur as the $f$-vector and Betti sequence of a finite simplicial complex, extending the well-known Euler–Poincaré theorem (cf. [6, 7]):
$\chi_{-1}=\sum_{j\geq 0}(-1)^{j}(f_{j}-\beta_{j})=1.$
Theorem 1.1 (Björner and Kalai, [2, Theorem 1.1]).
@@ 2. $(b)$
An ordered pair $(f,\beta)$ of sequences from $\mathbb{N}_{0}^{(\infty)}$ is called compatible, denoted by $f\sim\beta$, if there exists a simplicial complex with $f$-vector $f$ and Betti sequence $\beta$. By Theorem 1.1, this relation is purely combinatorial and independent of the field characteristic.
In their seminal work [2], Björner and Kalai further established the following foundational result concerning compatible pairs $(f,\beta)$.
Theorem 1.2 (Björner and Kalai, [2, Theorem 1.2, Theorem 5.3]).
@@ 6. $(f)$
We note that only the maximal part contains inequalities, which inspired us to introduce a new function to measure the discrepancy in these inequalities. The error function $\delta_{k-1}$ is defined by
$\delta_{k-1}(n)=\ell,$
where $\ell$ is the number of indices $t$ in the unique expansion (1) for which $a_{t}=t$. For completeness, we define $\delta_{k-1}(0)=0$.
In this paper, we prove the following result, which shows how the error function $\delta_{k}$ affects inequalities related to the combinatorial shadow functions $\partial_{k}$ and $\partial^{k}$.
Theorem 1.3.
Let $n$, $m$ be nonnegative integers and let $\epsilon_{k}$ be an integer with $0\leq\epsilon_{k}\leq\delta_{k}(n+m)$. Then the following statements hold:
  1. $(a)$
    $\partial_{k}(n)\leq m-\epsilon_{k}$ if and only if $n\leq\partial^{k}(n+m)$;
  2. $(b)$
    $n=\partial^{k}(n+m)$ if and only if $\partial_{k}(n)+\delta_{k}(n+m)=m$;
  3. $(c)$
    $\partial_{k}(n)=m$ if and only if $n=\partial^{k}(n+m)$ and $\delta_{k}(n+m)=0$.
As an application of Theorem 1.3, we sharpen the inequalities in parts $(c)$ and $(f)$ of Theorem 1.2 and provide an equivalent reformulation of part $(d)$. For a sequence $f=(f_{0},f_{1},\ldots)\in\mathbb{N}_{0}^{(\infty)}$, we define
@@ 6. $(f^{*})$
    $\phi(\psi(f)+\delta_{+}(f))=f-\delta(f)$ for all $f$-vectors $f$. Therefore, $\phi(\psi(f))\leq f-\delta(f)-\delta_{+}(f)$, with equality holding if and only if $\delta(f)=0$.
By applying Theorem 1.3 to Theorem 1.1, we also obtain the following result.
Corollary 1.5.
@@ 1. $(a)$
  2. $(b^{*})$
    $\chi_{-1}=1$, and $\partial_{k}(\chi_{k}+\beta_{k})\leq\chi_{k-1}-\epsilon_{k}$ for all $k\geq 1$,
  3. $(b^{+})$
    $\chi_{-1}=1$, and $0\leq f_{k}-\chi_{k-1}\leq\partial^{k}(f_{k})$ for all $k\geq 1$.
Remark 1.6.
It should be noted that condition $(b^{*})$ in Corollary 1.5 provides a refinement of condition $(b)$ from Theorem 1.1, while condition $(b^{+})$ establishes a new characterization of compatible pairs $(f,\beta)$. Furthermore, for any simplicial complex, the equality
$f_{k}-\chi_{k-1}=\dim Z_{k}$
holds, where $Z_{k}$ represents the space of $k$-cycles (see [2, Remark 4.4]). Therefore, through homological interpretation of the inequality $f_{k}-\chi_{k-1}\leq\partial^{k}(f_{k})$ in Corollary 1.5 $(b^{+})$, we obtain
$\dim Z_{k}\leq\partial^{k}(f_{k}).$
Let $m$ be a square-free positive integer, and let $P(m)$ be the set of prime factors of $m$. The family
$\Delta_{n}:=\{P(m):m\text{ is square-free and }m\leq n\}$
is a simplicial complex, called the Björner complex. It was shown by Björner [1] that the Prime Number Theorem and the Riemann Hypothesis are equivalent to certain asymptotic estimates of the reduced Euler characteristics of these complexes as $n\to\infty$.
Let $\sigma_{k}(n)$ (resp. $\sigma_{k}^{\mathrm{odd}}(n)$) denote the number of square-free integers (resp. odd square-free integers) $\leq n$ with exactly $k$ prime factors. In [1], Björner applied Theorem 1.1 to the Björner complex and proved that
$\partial_{k}(\sigma_{k+1}^{\mathrm{odd}}(n))\leq\sigma_{k}^{\mathrm{odd}}(n/2)\quad\text{for all }k\geq 1.$
Combining this result with Theorem 1.3, we immediately derive the following number-theoretic inequality.
Corollary 1.7.
For all $k\geq 1$,
$\sigma_{k+1}^{\mathrm{odd}}(n)\leq\partial^{k}(\sigma_{k+1}(n)).$
2 Proofs of Theorems 1.3 and 1.4
@@ 3. $(b^{+})$
such that $a_{k}>a_{k-1}>\cdots>a_{i}\geq i\geq 1$. This expansion is called the (k1)(k-1)-dimensional representation of $n$.
@@
are the $k$-dimensional representations of $n$ and $m$, respectively. Then the following statements hold:
  1. $(a)$
    $n=m$ if and only if $i=j$ and $a_{t}=b_{t}$ for all $t\in\{i,\ldots,k+1\}$;
  2. $(b)$
    $n<m$ if and only if either $i>j$ and $a_{t}=b_{t}$ for all $t\in\{i,\ldots,k+1\}$, or the maximal index $t\in\{\max\{i,j\},\ldots,k+1\}$ with $a_{t}\neq b_{t}$ satisfies $a_{t}<b_{t}$.
@@ 2. $(b)$
Note that $(a)$ follows from Lemma 2.1 immediately. Thus it suffices to consider $(b)$. Using Pascal’s identity $\binom{p}{q}=\binom{p-1}{q}+\binom{p-1}{q-1}$ recursively, we obtain
@@
for any integers $0\leq d\leq c+1$. This implies that
@@
whenever $c_{p}>d_{p}$ and $d_{p}>d_{p-1}>\cdots>d_{l}\geq l\geq 1$. Let $s=\max\{i,j\}$. If $a_{t}=b_{t}$ for all $t\in\{s,\ldots,k+1\}$, then $n<m$ when $i>j$, $n>m$ when $i<j$, and $n=m$ when $i=j$. Now suppose that there exists some index $t\in\{s,\ldots,k+1\}$ such that $a_{t}\neq b_{t}$. Let $t_{0}=\max\{t\in\{s,\ldots,k+1\}:a_{t}\neq b_{t}\}$. Then it follows from (2) that $n<m$ when $a_{t_{0}}<b_{t_{0}}$, and $n>m$ when $a_{t_{0}}>b_{t_{0}}$. This proves $(b)$. ∎
For any integer $n,k\geq 1$, using the unique expansion (1), we define:
$d^{k-1}(n)=\binom{a_{k}+1}{k}+\binom{a_{k-1}+1}{k-1}+\cdots+\binom{a_{i}+1}{i}$.
For completeness, define $d^{k-1}(0)=0$.
@@
The case $n=0$ is trivial, so we may assume $n\geq 1$. By Lemma 2.1, we suppose that the $k$-dimensional representation of $n$ is given by
@@
where $a_{k+1}>a_{k}>\cdots>a_{i}\geq i\geq 1$. Thus we have
@@
The case $n=0$ is trivial, so we may assume $n\geq 1$. By Lemma 2.1, we suppose that the $k$-dimensional representation of $n$ is given by
@@
where $a_{k+1}>a_{k}>\cdots>a_{i}\geq i\geq 1$. If $a_{t}=t$ for all $t\in\{i,\ldots,k+1\}$, then $\delta_{k}(n)=n=k+2-i$, $\partial^{k}(n)=\partial_{k}(\partial^{k}(n))=0$, and we are done. Now suppose that there exists some $t\in\{i,\ldots,k+1\}$ such that $a_{t}\neq t$. Let $t_{0}=\min\{t\in\{i,\ldots,k+1\}:a_{t}\neq t\}$. Then $a_{t}=t$ when $t<t_{0}$, and $a_{t}>t$ when $t\geq t_{0}$. According to the definition of the operator $\delta_{k}$, we obtain
@@
Therefore, by Lemma 2.3,
@@
Let $n$ and $m$ be nonnegative integers.
@@ 3. $(c)$
First consider $(a)$. If $d^{k}(n)=m$, by Lemma 2.2, we obtain $i=j$ and $a_{t}+1=b_{t}$ for all $t\in\{i,\ldots,k+1\}$. Therefore, $n=\partial^{k}(m)$. If $d^{k}(n)<m$, again by Lemma 2.2, either $i>j$ and $a_{t}+1=b_{t}$ for all $t\in\{i,\ldots,k+1\}$, or the maximal index $t\in\{\max\{i,j\},\ldots,k+1\}$ with $a_{t}+1\neq b_{t}$ satisfies $a_{t}+1<b_{t}$. For the former case, we have
@@
$n$ $=\binom{a_{k+1}}{k+1}+\binom{a_{k}}{k}+\cdots+\binom{a_{i}}{i}$
@@
Now consider $(b)$. If $n=m$, then $\partial_{k}(n)=\partial_{k}(m)$ by definition. Thus we can assume that $n<m$. Again by Lemma 2.2, we have that either $i>j$ and $a_{t}=b_{t}$ for all $t\in\{i,\ldots,k+1\}$, or the maximal index $t\in\{\max\{i,j\},\ldots,k+1\}$ with $a_{t}\neq b_{t}$ satisfies $a_{t}<b_{t}$. For the former case, we have
@@
For the latter case, let $t_{0}=\max\{t\in\{\max\{i,j\},\ldots,k+1\}:a_{t}\neq b_{t}\}$. Then $a_{t_{0}}<b_{t_{0}}$, and we obtain
@@
If $\partial_{k}(n)\leq m-\epsilon_{k}$, by Lemma 2.3,
$d^{k}(n)=n+\partial_{k}(n)\leq n+m-\epsilon_{k}.$
@@
$n\leq\partial^{k}(n+m-\epsilon_{k})=\partial^{k}(n+m).$
Conversely, if $n\leq\partial^{k}(n+m)$, by Lemma 2.5 $(b)$, we have
$\partial_{k}(n)\leq\partial_{k}(\partial^{k}(n+m)),$
@@
$n+\partial_{k}(n)\leq\partial^{k}(n+m)+\partial_{k}(\partial^{k}(n+m)).$
@@
$\partial_{k}(n)$ $\leq m-(n+m-\partial^{k}(n+m)-\partial_{k}(\partial^{k}(n+m)))=m-\delta_{k}(n+m)\leq m-\epsilon_{k}.$
This proves $(a)$.
If $n=\partial^{k}(n+m)$, then $\partial_{k}(n)=\partial_{k}(\partial^{k}(n+m))$. This implies
$n+\partial_{k}(n)=\partial^{k}(n+m)+\partial_{k}(\partial^{k}(n+m)),$ (3)
@@
$\partial_{k}(n)+\delta_{k}(n+m)=m$
by Lemma 2.4. Conversely, if $\partial_{k}(n)+\delta_{k}(n+m)=m$, then we recover (3). By Lemma 2.3, it follows that
$d^{k}(n)=d^{k}(\partial^{k}(n+m)).$
@@
Suppose that $n=\partial^{k}(n+m)$. According to $(b)$, we have $\partial_{k}(n)+\delta_{k}(n+m)=m$. If $\delta_{k}(n+m)=0$ then $\partial_{k}(n)=m$, and if $\delta_{k}(n+m)\neq 0$, then $\partial_{k}(n)\neq m$. On the other hand, if $\partial_{k}(n)=m$, by Lemma 2.3,
$d^{k}(n)=\partial_{k}(n)+n=m+n.$
Combining this with Lemma 2.5 $(a)$, we obtain $n=\partial^{k}(n+m)$, and so $(c)$ follows.
@@
In order to prove Theorem 1.4, we also need the Kruskal-Katona theorem [4, 5], which gives a complete characterization of $f$-vectors of simplicial complexes.
@@
First observe that $(d^{*})$ follows from Theorem 1.2 and Theorem 1.3 immediately. Now suppose that $f$ is the $f$-vector of some simplicial complex. Let $\beta\in B_{f}$. Since $f\sim\beta$, by Corollary 1.5, we have $\chi_{-1}=1$, and $f_{k}-\chi_{k-1}\leq\partial^{k}(f_{k})$ for all $k\geq 1$. Also note that $f_{0}-\chi_{-1}=f_{0}-1=\partial^{0}(f_{0})$. Hence, for each $k\geq 0$,
@@
Furthermore, according to [2, Theorem 5.3], the upper bound in (4) can be achieved uniformly, that is,
@@
Next we consider $(f^{*})$. According to $(c^{*})$ and $(d^{*})$, if $f$ is an $f$-vector with $\delta(f)=0$, then $\phi(\psi(f))=f$. Now suppose that $f$ is an arbitrary $f$-vector. By Lemma 2.6, it is easy to see that $f-\delta(f)$ is also an $f$-vector. Moreover, by definition, we have $\delta(f-\delta(f))=0$. Thus $\phi(\psi(f-\delta(f)))=f-\delta(f)$ by the above arguments. Note that, for any $k\geq 0$,
@@
By [2, Theorem 5.2], the mapping $\phi$ is injective and order-preserving. Therefore, we conclude that
@@
where equalities hold everywhere if and only if $\delta(f)=0$.